A group of 3 children and 2 adults pay
a total of $120 to take a karate class.
A group of 5 children and 1 adult take
the same karate class for $95. What is
the total cost for 1 child and 1 adult to
take the karate class?

Answer by richwmiller(17219)

child=c
a=adult
3c+2a=120;
5c+1a=95
a=95-5c
3c+2*(95-5c)=120
3c-10c+190=120
-7c=-70
c=10
30+2a=120
2a=90
a=45